# Arena: A Patch-of-Interest ViT Inference Acceleration System for Edge-Assisted Video Analytics

### v1

@@ Abstract.
The advent of edge computing has made real-time intelligent video analytics feasible. Previous works, based on traditional model architecture (e.g., CNN, RNN, etc.), employ various strategies to filter out non-region-of-interest content to minimize bandwidth and computation consumption but show inferior performance in adverse environments. Recently, visual foundation models based on transformers have shown great performance in adverse environments due to their amazing generalization capability. However, they require a large amount of computation power, which limits their applications in real-time intelligent video analytics. In this paper, we find visual foundation models like Vision Transformer (ViT) also have a dedicated acceleration mechanism for video analytics. To this end, we introduce Arena, an end-to-end edge-assisted video inference acceleration system based on ViT. We leverage the capability of ViT that can be accelerated through token pruning by only offloading and feeding Patches-of-Interest (PoIs) to the downstream models. Additionally, we employ probability-based patch sampling, which provides a simple but efficient mechanism for determining PoIs where the probable locations of objects are in subsequent frames. Through extensive evaluations on public datasets, our findings reveal that Arena can boost inference speeds by up to 1.58$\times$ and 1.82$\times$ on average while consuming only 54% and 34% of the bandwidth, respectively, all with high inference accuracy.
Video Analytics System, Edge Computing, Vision Transformer
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; October 28 - November 1, 2024; Melbourne, Australia.isbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Computer vision tasksccs: Computer systems organization Real-time system architecture
1. Introduction
@@
Today, IoT cameras are generating an unprecedented volume of video data daily. According to predictions by Hailo (Hailo, 2024), the global market is anticipated to see approximately 180 million shipments of embedded video surveillance devices by 2025. These cameras power a plethora of video analytics applications such as traffic management (Tong et al., 2021; Li et al., 2023), autonomous driving (Shi et al., 2022; Zhang et al., 2021b), security surveillance (Yang et al., 2023; Yi et al., 2020), etc.
@@
The massive data generated by IoT cameras far exceeds their limited computing capabilities. A straightforward way is to offload the captured data to edge servers. These servers execute intensive downstream inference tasks, usually based on deep neural networks, and deliver useful visual feedback to users. Previous works have explored the spatial and temporal redundancy of video content to design systems that reduce the bandwidth usage between cameras and edge servers. Efforts are made either to extract and transmit regions of interest (RoIs) from video frames (Liu et al., 2022b; Zhang et al., 2021a) or to design a module that filters out similar frames (Tchaye-Kondi et al., 2022; Chen et al., 2015). Though these methods have optimized the transmission and computation, their downstream models based on Convolutional Neural Network (CNN) show inferior performance in the live video data, because of the poor generalization ability of CNN.
@@
Figure 1. Arena: our patch-of-interest ViT inference acceleration system for edge-assisted video analytics. Due to the limited computing power of the camera, the extracted patches-of-interest are offloaded to an edge server for processing with its more powerful GPUs. MTPs stands for Memory Token Pools.
[…]
Our Edge-assisted computation offloading in video analytics.
@@
In recent years, visual foundation models based on transformer architecture (Dosovitskiy et al., 2020; Liu et al., 2022a, 2021; Li et al., 2022), exemplified by Vision Transformer (ViT) (Dosovitskiy et al., 2020), have surpassed CNN-based methods (Girshick, 2015; Redmon and Farhadi, 2018; Cai and Vasconcelos, 2019; Tian et al., 2019) in most downstream tasks, including image classification (Deng et al., 2009), object detection (Lin et al., 2014) and object tracking (Milan et al., 2016). The strengths of the ViT lie in its exceptional feature extraction capabilities and generalization. Through its multi-head self-attention mechanism (MSA) (Vaswani et al., 2017), ViT can more effectively capture global information within the input, achieving superior performance in various aspects. Furthermore, the ”pre-training + fine-tuning” paradigm has also standardized the implementation process of video analytics, enabling models to be rapidly deployed in entirely new specific scenes using only a few samples. However, the computational demands of its MSA increase quadratically with the number of patches. This intense computational requirement, especially in high-resolution video (e.g., 1080p) analytics scenarios, impedes the practical application deployment of ViT-based models. Although there exits works on efficient ViT that utilize token pruning (Rao et al., 2021; Liang et al., 2022), merging (Bolya et al., 2022; Zeng et al., 2022), and fusion (Kim et al., 2024) techniques to reduce computation, these approaches still require a full image as input and fail to adapt to real-time intelligent video analytics.
@@
To fill this gap, we propose Arena, a Patch-of-Interest (PoI) ViT inference acceleration system for edge-assisted video analytics. We find that ViT also has a dedicated acceleration mechanism for video analytics. As shown in Figure 1, first, the camera transmits a full video frame (referred to as a keyframe) to the edge server, followed by conducting inference on it, where we establish and maintain memory token pools. Subsequently, we utilize a Probability-based Patch Sampling (PPS) mechanism to determine the PoIs for the next frame (referred to as a Non-keyframe). Therefore, only PoIs, usually occupying a small portion of the frame, are required to transmit to the edge server. Then, a model based on the ViT backbone performs inference merely using these PoIs. Finally, a Memory Feature Reconstruction (MFR) module is employed to restore the complete feature maps, making it available for a dense prediction task head. Arena enhances overall system throughput and computational efficiency twofold. On the one hand, transmitting only partial patches significantly reduces bandwidth consumption and transmission latency. On the other hand, feeding PoIs to the ViT backbone and reconstructing features from memory tokens reduce computational overhead and inference time while almost maintaining the same accuracy.
@@
We develop and deploy a prototype on a testbed that uses an NVIDIA Jetson as a camera device and an Ubuntu desktop as an edge server running real video analytics workloads. Our results show that Arena can accelerate the inference up to 1.58×1.58$\times$1.58 × and 1.82×1.82$\times$1.82 × on average by only using 54% and 34% bandwidth while keeping the inference accuracy sacrifice within an acceptance range.
@@
The remainder of this paper is organized as follows. We present the preliminaries and motivation in Section 2. The design of Arena is introduced in Section 3, followed by the system implementation in Section 4. We conduct extensive experiments in Section 5, and the related work is reviewed in Section 6. Finally, we conclude this paper in Section 7.
2. Preliminaries And Motivation
@@
We first introduce the preliminaries of Vision Transformer (Sec. 2.1). Next, we reveal that video data contains a significant amount of redundancy. Substantial bandwidth can be saved by only transmitting the PoIs (Sec. 2.2). ViT is well-suited for unstructured data because it can process variable-length sequences as inputs. This capability allows for filtering out data at the beginning of the pipeline and accelerates the computation (Sec. 2.3), which gives us insights on how to accelerate ViT-based video analytics systems.
2.1. Preliminaries of Vision Transformer
@@
(1) $\tilde{z}_{0}=\left[x_{class},x_{p,1}\mathbf{E};x_{p,2}\mathbf{E};\ldots;x_{p,N}\mathbf{E}\right],$
@@
(2) $z_{0}=\tilde{z}_{0}+\mathbf{E}_{pos},\qquad\mathbf{E}_{pos}\in\mathbb{R}^{(N+1)\times D}.$
In the end, the transformer encoder (Vaswani et al., 2017) consisting of MSA, MLP blocks, and LayerNorm (LN) is adopted to update $z_{0}$ with $L$ iteration and output the final image representation $\mathbf{y}$. The formulations are
(3) $z_{\ell}^{\prime}$ $=\operatorname{MSA}\left(\operatorname{LN}(z_{\ell-1})\right)+z_{\ell-1}^{\prime},$ $\ell=1\ldots L,$
@@
(4) $\mathbf{y}=\mathrm{LN}(\mathbf{z}_{L}^{0}).$
2.2. Redundancy in Video Inference Data
In edge-assisted video analytics, high-resolution videos often include much redundant data. Particularly, the video frame is divided into multiple patches of certain pixels (e.g., $16\times 16$), among which those containing the objects are referred to as PoIs. Conversely, the rest of the frame is dominated by the background (e.g., buildings and sky) and other irrelevant objects (Jiang et al., 2021). Table 1 shows the redundancy of the first five cameras from two popular real-world video datasets. It is evident that in most videos, the proportion of PoI does not exceed 25%. Additionally, in some specific scenes, PoIs constitute no more than 5% of the whole frame. By only transmitting the PoIs from the camera to the edge server, most videos can achieve a bandwidth usage reduction exceeding 80%, with potential savings reaching up to 92.01%. This discovery motivates us to seek a more fine-grained method for extracting PoIs that allows transmitting as few patches as possible while still meeting the accuracy requirements.
Table 1. Redundancy in video inference data on MOT17Det (Milan et al., 2016) and AIC22 (Naphade et al., 2022) datasets.
@@ 2.2. Redundancy in Video Inference Data
2.3. Accelerating ViT via Patch Prunning
Figure 2. Pruning patches can accelerate ViT backbone inference. We evaluate the impact of pruning 25%, 50%, and 75% tokens on (a) inference latency and (b) GFlops using ViT-base across videos of 1080p, 720p, and 480p resolutions. Similar trends were found in ViT-Base.
One of the desirable advantages of ViT is its ability to allow a flexible length of patch input. Specifically, the total computational complexity of one MSA module and one MLP is $\mathcal{O}\left(12ND^{2}+2N^{2}D\right)$. Obviously, reducing the number of patches can linearly or even quadratically reduce the operations. Figure 2 illustrates the (a) average inference latency and (b) average GFlops per frame when different proportions of patches are pruned across videos of various resolutions. It is noteworthy that using ViT-small on 1080p videos and pruning 75% patches resulted in a reduction of latency by 62ms, only 11.76% of the original latency. Additionally, the GFlops are linearly related to the number of input patches, showing a reduction of approximately 74.97% if pruning 75% patches. Furthermore, we confirm this on video datasets in Table 1. The Latency column reports the average inference latency saved by only feeding the PoIs to the ViT backbone for each frame. Most videos can achieve over an 80% acceleration, with the highest reaching up to 90.11%. This observation shows the potential opportunities for accelerating ViT inference by pruning the number of input patches.
@@ 2.3. Accelerating ViT via Patch Prunning
3. Arena Design
@@
Figure 3. Given $K$ continuous frames $\{\hat{\mathbf{x}}^{1},\mathbf{x}^{2},\ldots,\mathbf{x}^{K}\}$ in an interval, Arena periodically operates in two distinct phases: keyframe inference (Left) for the first frame $\hat{\mathbf{x}}^{1}$ and non-keyframe inference (Right) for the rest of the frames. Both two phases use the same network architecture with shared weights. Notably, we split the frame into nine patches only for demonstration.
3.1. Overview
@@
Arena is an edge-assisted, real-time inference acceleration system tailored for ViT-based models. It aims to reduce communication and computation overheads by pruning the number of patches offloaded to the server while maintaining high accuracy. Arena comprises a camera and an edge server. The camera, equipped with limited computational power, continuously captures video frames, and the edge server deploys a ViT-based detector to perform inference. In every $K$ frames, the first frame is designated as the keyframe, and the subsequent $K-1$ frames are considered as non-keyframes. Arena operates in two distinct phases: keyframe inference (Sec. 3.2) and non-keyframe inference (Sec. 3.3), then alternating periodically between them as illustrated in Figure 3. In keyframe inference, Arena first performs full frame inference by detector and caches intermediate tokens into two memory token pools. In the non-keyframe inference phase, it queries memory tokens from the pools and reconstructs complete image-wise features to accommodate unstructured samples with variable-length patch sequences. Moreover, we introduce PPS to identify the PoIs in non-keyframes and provide feedback to the camera (Sec. 3.4). Finally, Sec. 3.5 describes our training approach in a video-training manner. Notably, both phases employ the same network and shared weights. We take a sequence of frames $\{\hat{\mathbf{x}}^{1},\mathbf{x}^{2},\ldots,\mathbf{x}^{K}\}$ as an example, where $\hat{\mathbf{x}}^{1}$ denotes the keyframe and $\mathbf{x}^{2},\ldots,\mathbf{x}^{K}$ as non-keyframes, and detail our above-mentioned mechanisms in the following paragraphs.
3.2. Keyframe Inference

### v2

# Patch-of-Interest ViT Inference Acceleration System for Edge-Assisted Video Analytics

@@ Abstract
The advent of edge computing has made real-time intelligent video analytics feasible. Previous works, based on traditional model architecture (e.g., CNN, RNN, etc.), employ various strategies to filter out non-region-of-interest content to minimize bandwidth and computation consumption but show inferior performance in adverse environments. Recently, visual foundation models based on transformers have shown great performance in adverse environments due to their amazing generalization capability. However, they require a large amount of computation power, which limits their applications in real-time intelligent video analytics. In this paper, we find visual foundation models like Vision Transformer (ViT) also have a dedicated acceleration mechanism for video analytics. To this end, we introduce Arena, an end-to-end edge-assisted video inference acceleration system based on ViT. We leverage the capability of ViT that can be accelerated through token pruning by only offloading and feeding Patches-of-Interest to the downstream models. Additionally, we design an adaptive keyframe inference switching algorithm tailored to different videos, capable of adapting to the current video content to jointly optimize accuracy and bandwidth. Through extensive experiments, our findings reveal that Arena can boost inference speeds by up to 1.58$\times$ and 1.82$\times$ on average while consuming only 47% and 31% of the bandwidth, respectively, all with high inference accuracy.
Index Terms:
Video Analytics System, Edge Computing, Vision Transformer.
I Introduction
@@
Today, IoT cameras are generating an unprecedented volume of video data daily. According to predictions by Hailo [1], the global market is anticipated to see approximately 180 million shipments of embedded video surveillance devices by 2025. These cameras power a plethora of video analytics applications such as traffic management [2], autonomous driving [3], security surveillance [4], etc.
@@
The massive data generated by IoT cameras far exceeds their limited computing capabilities. A straightforward way is to offload the captured data to edge servers [5, 6, 7, 8, 9, 10, 11, 12, 13, 14]. These servers execute intensive downstream inference tasks, usually based on deep neural networks, and deliver useful visual feedback to users. Previous works have explored the spatial and temporal redundancy of video content to design systems that reduce the bandwidth usage between cameras and edge servers. Efforts are made either to extract and transmit regions of interest (RoIs) from video frames [8, 10] or to design a module that filters out similar frames [11, 15]. Though these methods have optimized the transmission and computation, their downstream models based on Convolutional Neural Network (CNN) show inferior performance in the live video data, because of the poor generalization ability of CNN.
@@
Figure 1: Arena: our patch-of-interest ViT inference acceleration system for edge-assisted video analytics. Due to the limited computing power of the camera, the extracted patches-of-interest are offloaded to an edge server for processing with its more powerful GPUs. MTPs stands for Memory Token Pools.
@@
In recent years, visual foundation models based on transformer architecture [16, 17, 18], exemplified by Vision Transformer (ViT) [16], have surpassed CNN-based methods in most downstream tasks, including image classification, object detection and object tracking. The strengths of the ViT lie in its exceptional feature extraction capabilities and generalization. Through its multi-head self-attention mechanism (MSA), ViT can more effectively capture global information within the input, achieving superior performance in various aspects. Furthermore, the “pre-training + fine-tuning” paradigm has also standardized the implementation process of video analytics, enabling models to be rapidly deployed in entirely new specific scenes using only a few samples. However, the computational demands of its MSA increase quadratically with the number of patches. This intense computational requirement, especially in high-resolution video (e.g., 1080p) analytics scenarios, impedes the practical application deployment of ViT-based models. Although there exits works on efficient ViT that utilize token pruning [19, 20], merging [21, 22], and fusion [23] techniques to reduce computation, these approaches still require a full image as input and fail to adapt to real-time intelligent video analytics.
@@
To fill this gap, we propose Arena, a Patch-of-Interest (PoI) ViT inference acceleration system for edge-assisted video analytics. We find that ViT also has a dedicated acceleration mechanism for video analytics. As shown in Figure 1, first, the camera transmits a full video frame (referred to as a keyframe) to the edge server, followed by conducting inference on it, where we establish and maintain memory token pools. Subsequently, we utilize a Probability-based Patch Sampling (PPS) mechanism to determine the PoIs for the next few frames (referred to as a Non-keyframe). Therefore, only PoIs, usually occupying a small portion of the frame, are required to transmit to the edge server. Then, a model based on the ViT backbone performs inference merely using these PoIs. Thus, inference on non-keyframes significantly reduces the workload compared to keyframe inference. Finally, a Memory Feature Reconstruction (MFR) module is employed to restore the complete feature maps, making it available for a dense prediction task head.
Arena periodically alternates between the keyframe inference and non-keyframe inference phases. Due to the temporal and spatial characteristics of different cameras, the amount and timing of the appearance of task-relevant objects vary as described in [24]. For example, on a street at night, the number of interesting objects, such as people and cars, is much fewer compared to daytime. Similarly, the number of objects in remote rural videos differs significantly from that in urban areas. Therefore, a fixed strategy for switching between keyframes and non-keyframes is not suitable for all video analytics scenarios. Consequently, we propose an Adaptive Keyframe Interval Switching (AKIS) algorithm to control when to switch between keyframe inference and non-keyframe inference.
Arena enhances overall system throughput and computational efficiency twofold. On the one hand, transmitting only partial patches significantly reduces bandwidth consumption and transmission latency. On the other hand, feeding PoIs to the ViT backbone and reconstructing features from memory tokens reduce computational overhead and inference time while maintaining almost the same accuracy. Additionally, the AKIS algorithm can personalize the configuration of the keyframe interval for each video, further achieving a trade-off between accuracy and bandwidth. We develop and deploy a prototype on a testbed that uses an NVIDIA Jetson as a camera device and an Ubuntu desktop as an edge server running real video analytics workloads. We summarize the contributions of this work as follows:
  • We propose Arena, an edge-assisted video analytics system that utilizes PoI ViT inference, accelerating inference while reducing network bandwidth.
  • We have designed an adaptive keyframe interval switching algorithm that adapts to the current video content, jointly optimizing accuracy and bandwidth in real-time.
@@
  • Experimental results on the testbed show that Arena can accelerate the inference up to 1.58×1.58$\times$1.58 × and 1.82×1.82$\times$1.82 × on average by only using 47% and 31% bandwidth while keeping the inference accuracy sacrifice within an acceptance range.
@@
The remainder of this paper is organized as follows. We present the preliminaries and motivation in Section II. The design of Arena is introduced in Section III, followed by the system implementation in Section IV. We conduct extensive experiments in Section V, and the related work is reviewed in Section VI. Finally, we conclude this paper in Section VII.
II Preliminaries And Motivation
@@
We first introduce the preliminaries of Vision Transformer (Sec. II-A). Next, we reveal that video data contains a significant amount of redundancy. Substantial bandwidth can be saved by only transmitting the PoIs (Sec. II-B). However, traditional CNN-based models cannot efficiently process these unstructured regions (Sec. II-C). In contrast, ViT is well-suited for unstructured data because it can process variable-length sequences as inputs. This capability allows for filtering out data at the beginning of the pipeline and accelerates the computation (Sec. II-D), which gives us insights on how to accelerate ViT-based video analytics systems. Finally, the varying trade-offs induced by keyframe intervals across different scenarios (Sec. II-E) inspire us to develop methods for keyframe interval adjustment.
II-A Preliminaries of Vision Transformer
@@
$\tilde{z}_{0}=\left[x_{class},x_{p,1}\mathbf{E};x_{p,2}\mathbf{E};\ldots;x_{p,N}\mathbf{E}\right],$ (1)
@@
$z_{0}=\tilde{z}_{0}+\mathbf{E}_{pos},\qquad\mathbf{E}_{pos}\in\mathbb{R}^{(N+1)\times D}.$ (2)
In the end, the transformer encoder consisting of MSA, MLP blocks, and LayerNorm (LN) is adopted to update $z_{0}$ with $L$ iteration and output the final image representation $\mathbf{y}$. The formulations are
$z_{\ell}^{\prime}$ $=\operatorname{MSA}\left(\operatorname{LN}(z_{\ell-1})\right)+z_{\ell-1}^{\prime},$ $\ell=1\ldots L,$ (3)
@@
$\mathbf{y}=\mathrm{LN}(z_{L}[0]).$ (4)
II-B Redundancy in Video Inference Data
In edge-assisted video analytics, high-resolution videos often include much redundant data. Particularly, the video frame is divided into multiple patches of certain pixels (e.g., $16\times 16$), among which those containing the objects are referred to as PoIs. Conversely, the rest of the frame is dominated by the background (e.g., buildings and sky) and other irrelevant objects [25]. Table I shows the redundancy of the first five cameras from two popular real-world video datasets. It is evident that in most videos, the proportion of PoI does not exceed 25%. Additionally, in some specific scenes, PoIs constitute no more than 5% of the whole frame. By only transmitting the PoIs from the camera to the edge server, most videos can achieve a bandwidth usage reduction exceeding 80%, with potential savings reaching up to 92.01%. This discovery motivates us to seek a more fine-grained method for extracting PoIs that allows transmitting as few patches as possible while still meeting the accuracy requirements.
TABLE I: Redundancy in video inference data on MOT17Det [26] and AIC22 [27] datasets.
@@ II-B Redundancy in Video Inference Data
II-C The Limitation of CNN-based Backbone on Unstructured Input
Figure 2: The inference latency for three strategies: Full Frame, Masked Frame, and RoIs separately. Downstream models based on CNNs fail to benefit from the filtered RoIs.
Models based on a CNN backbone struggle to handle unstructured inputs. Previous approaches use lightweight methods to filter RoI from the video frame [5, 8, 28, 29], which is indeed effective in reducing bandwidth consumption. However, they either infer masked, full-size images or process each RoI individually. The former does not reduce the computation on the backbone because of the same size, while the latter requires multiple inferences and cannot batch together due to varying RoI sizes. For example, we follow the strategy from [5] to extract these regions whose objectness scores are greater than 50%, as shown in Figure 2a. Figures 2b and 2c illustrate the masked image and RoI respectively. We employ a full detector to infer these two types of unstructured data, with latencies of 73.215ms and 1855.122ms, respectively, which do not achieve acceleration compared to the original image’s inference time of 74.077ms. Therefore, downstream models based on CNNs fail to achieve acceleration benefits from the filtered RoIs.
II-D Accelerating ViT via Patch Prunning
Figure 3: Pruning patches can accelerate ViT backbone inference. We evaluate the impact of pruning 25%, 50%, and 75% tokens on (a) inference latency and (b) GFlops using ViT-base across videos of 1080p, 720p, and 480p resolutions. Similar trends were found in ViT-Base.
One of the desirable advantages of ViT is its ability to allow a flexible length of patch input. Specifically, the total computational complexity of one MSA module and one MLP is $\mathcal{O}\left(12ND^{2}+2N^{2}D\right)$. Obviously, reducing the number of patches can linearly or even quadratically reduce the operations. Figure 3 illustrates the (a) average inference latency and (b) average GFlops per frame when different proportions of patches are pruned across videos of various resolutions. It is noteworthy that using ViT-small on 1080p videos and pruning 75% patches resulted in a reduction of latency by 62ms, only 11.76% of the original latency. Additionally, the GFlops are linearly related to the number of input patches, showing a reduction of approximately 74.97% if pruning 75% patches. Furthermore, we confirm this on video datasets in Table I. The Latency column reports the average inference latency saved by only feeding the PoIs to the ViT backbone for each frame. Most videos can achieve over an 80% acceleration, with the highest reaching up to 90.11%. This observation shows the potential opportunities for accelerating ViT inference by pruning the number of input patches.
@@ II-D Accelerating ViT via Patch Prunning
Figure 4: Keyframe interval introduces accuracy and bandwidth trade-off in different scenes.
II-E Keyframe Interval Introduces Accuracy and Bandwidth Trade-off
Cameras are commonly deployed in dynamic scenes such as traffic intersections, building entrances, and pedestrian streets, where the quantity of PoIs changes frequently. A naive keyframe scheduling policy picks a keyframe at a pre-fixed rate (e.g., every $K$ frame [30]) is not feasible for different scenes.
We utilize two distinct types of video datasets and apply fixed keyframe interval strategies with varying intervals. MOT17Det is a dataset containing pedestrians on moving or stationary streets, where the number of objects in each frame is small, and their movement speed is slow. In contrast, AIC22 is a traffic dataset comprising moving vehicles, with a high number of objects moving at fast speeds. As illustrated in Fig. 4, increasing the keyframe interval in MOT17Det has a negligible impact on accuracy, while significantly reducing bandwidth consumption. However, in AIC22, although increasing the keyframe interval similarly reduces bandwidth consumption, it also leads to a linear decrease in accuracy. Different keyframe intervals thus present distinct trade-offs across various datasets. This phenomenon is observed not only across different datasets with varying objects but can also occur across different videos in the same dataset, as the moving speed of objects in videos is constantly changing.
The above experiments demonstrate that a fixed keyframe interval is not suitable for videos with dynamic changes. Designing a strategy that adaptively adjusts the keyframe interval to balance accuracy and bandwidth in a real-time video analytics system is a crucial challenge for Arena.
III Arena Design
@@
Figure 5: The overview of Arena. Given $K$ continuous frames $\{\hat{\mathbf{x}}^{1},\mathbf{x}^{2},\ldots,\mathbf{x}^{K}\}$ in an interval, Arena periodically operates in two distinct phases: keyframe inference (Left) for the first frame $\hat{\mathbf{x}}^{1}$ and non-keyframe inference (Right) for the rest of the frames. AKIS (Down), deployed on the camera, utilizes information from historical frames to determine subsequent keyframe intervals. Notably, we split the frame into nine patches only for demonstration.
III-A Overview
@@
Arena is an edge-assisted, real-time inference acceleration system tailored for ViT-based models. It aims to reduce communication and computation overheads by pruning the number of patches offloaded to the server while maintaining high accuracy. Arena comprises a camera and an edge server. The camera, equipped with limited computational power, continuously captures video frames, and the edge server deploys a ViT-based detector to perform inference. In every $K$ frames, the first frame is designated as the keyframe, and the subsequent $K-1$ frames are considered as non-keyframes. Arena operates in two distinct phases: keyframe inference (Sec. III-B) and non-keyframe inference (Sec. III-C), then alternating periodically between them as illustrated in Figure 5. In keyframe inference, Arena first performs full frame inference by detector and caches intermediate tokens into two memory token pools. In the non-keyframe inference phase, it queries memory tokens from the pools and reconstructs complete image-wise features to accommodate unstructured samples with variable-length patch sequences. Moreover, we introduce PPS to identify the PoIs in non-keyframes and provide feedback to the camera (Sec. III-D). Additionally, the AKIS algorithm is proposed to personalize the settings for keyframe and non-keyframe intervals, further optimizing the trade-off between accuracy and bandwidth (Sec. III-E). Finally, Sec. III-F describes our training approach in a video-training manner. Notably, both phases employ the same network and shared weights. We take a sequence of frames $\{\hat{\mathbf{x}}^{1},\mathbf{x}^{2},\ldots,\mathbf{x}^{K}\}$ as an example, where $\hat{\mathbf{x}}^{1}$ denotes the keyframe and $\mathbf{x}^{2},\ldots,\mathbf{x}^{K}$ as non-keyframes, and detail our above-mentioned mechanisms in the following paragraphs.
III-B Keyframe Inference